OpenStudy (anonymous):

find the volume of a regular tetrahedron having a side of 20 cm.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

942.81 cm

OpenStudy (anonymous):

how? :)

OpenStudy (anonymous):

easy type your question on the internet. and the tell me

OpenStudy (anonymous):

wolf am i right

OpenStudy (nirmalnema):

V=a^3 / 6rt2

OpenStudy (wolf1728):

gee I just got here. I'd need to look up the formula for a regular tetrahedron. seems nirma has it.

OpenStudy (anonymous):

tnx :)

OpenStudy (wolf1728):

Volume = edge^3 / (6*sqrt(2) Volume = 20^3 / 8.4852813742 Volume = 8,000 / 8.4852813742 Volume = 942.8090415821 Volume = 942.81 rounded and that agrees with 72595210002's calculation